10. What is the frequency of the A allele in the United States?








11. In a population of 330,000,000 people, how many do we expect to be carriers for albinsim (but not show the albino phenotype)?






Coat color in dogs is determined by one gene with two alleles, B (brown fur) and b (yellow fur); simple dominance rules apply. In a population of 500 Labrador retrievers, you see the following genotypes:

BB​(Brown Fur)​180 individuals​

Bb​(Brown Fur)​240 individuals ​

bb​(Yellow Fur)​80 individuals ​​

*** For #1- 4 Do NOT assume population is in Hardy-Weinberg Equilibrium ***

1. Calculate the allele frequencies of B.

2. Calculate the allele frequencies of b.

In the following generation of Labrador retrievers, the genotypes are observed to be:

BB​(Brown Fur)​234 individuals​

Bb​(Brown Fur)​312 individuals ​

​bb​(Yellow Fur)​104 individuals

3. Calculate the allele frequencies of B.

4. Calculate the allele frequencies of b.

5. Is the population of Labradors above in Hardy-Weinberg Equilibrium?

In a species of lizard, a single gene determines scale color. The gene has two alleles, G (green) and g (blue); simple dominance rules apply. In a sampled population, it is determined that the g allele has a frequency of 0.24.

6. What is the frequency of the G allele?

7. What are the three possible genotypes and their corresponding frequencies?

8. In a population of 5,000 lizards, how many would you expect to have blue scales(bb)?

Albinism is a genetic disorder where there is little to no production of the pigment melanin in the skin, hair or eyes. The albino phenotype is only expressed in homozygous recessive individuals. In the United States, the frequency of albinism is 1 in 20,000. You can use A to represent the dominant, non-albinism allele and a to represent the recessive allele.

9. What is the frequency of the a allele in the United States?

10. What is the frequency of the A allele in the United States?

11. In a population of 330,000,000 people, how many do we expect to be carriers for albinsim (but not show the albino phenotype)?

Albinism occurs at different frequencies across the animal kingdom. In birds, albinism is estimated to occur in 1 in 1,764. In mammals (excluding humans), albinism is estimated to occur in 1 in 10,000.

12. What is the frequency of the a allele in birds?

13. What is the frequency of the a allele in mammals?

In polar bears, body size directly impacts fitness through males’ ability to win competitions and females’ ability to provide for offspring. We want to determine whether body size, in polar bears, is a heritable trait. A survey of 756 polar bears found the average body size to be 250 cm.

A selected group of the 30 largest bears (average body size is 280cm) are later bred and the body size of their offspring (at maturity) is measured. The average body size of their offspring is 260 cm.

14. Calculate the selection differential.

15. Calculate the response to selection.

16. Calculate heritability (h2).

17. Does this trait appear to be strongly influenced by genetics, environment, or a combination? Explain.




Sorry I posted #10 and #11 twice
Answer #1

Answer:

1).

BB 180 = 180*2 = 360 B alleles

Bb 240 = 240 B * 240 b alleles

bb 80 = 160 b alleles

Total alleles = 1000

Total B alleles = 360+240 = 600

Frequency of B allele = 600/1000 = 0.6

2).

Total b alleles = 160+240 = 400

Frequency of b allele = 400/1000 = 0.4

3).

BB 234 = 234*2 = 468 B alleles

Bb 312 = 312 B * 312 b alleles

bb 104 = 208 b alleles

Total alleles = 1300

Total B alleles = 468+312 = 780

Frequency of B allele = 780/1300 = 0.6

4).

Total b alleles = 312+208 = 520

Frequency of b allele = 520/1300 = 0.4

5).

Hardy-weinberg equilibrium = p+q=1

0.6+0.4 = 1

Hence, the population is in Hardy-Weinberg equilibrium

6). The frequency of the G allele = 1-0.24 = 0.76 (asp+q=1)

7).

GG = 0.76*0.76 = 0.58

Gg = 2*0.76*0.24 = 0.36

gg = 0.24*0.24 = 0.06

8).

Number of blue individuals = 0.06*5000 = 300